# Learning-based Axial Motion Magnification

### v1

@@ 3.3 Neural Networks and Training
Figure 3 shows the synthetic data generation pipeline. We sample one background from COCO [18] and $K{-}1$ number of foreground textures with segmentation masks from PASCAL VOC [10]. These elements are randomly located on image planes of resolution 384$\times$384 to produce $K$ previous layer images $\{{\mathbf{L}}^{k}_{1}\}_{k=1}^{K}$ and corresponding masks $\{\boldsymbol{\Omega}^{k}_{1}\}_{k=1}^{K}$. Following this, with randomly sampled $K$ translation parameters $\{\textbf{d}^{k}\}_{k=1}^{K}$, we generate the next layer images $\{{\mathbf{L}}^{k}_{2}\}_{k=1}^{K}$ and masks $\{\boldsymbol{\Omega}^{k}_{2}\}_{k=1}^{K}$ by translating the initial layers and masks according to $\{\textbf{d}^{k}\}_{k=1}^{K}$. For the axially magnified layer images $\{\hat{{\mathbf{L}}}^{\phi,k}\}_{k=1}^{K}$ and their masks $\{\hat{\boldsymbol{\Omega}}^{\phi,k}\}_{k=1}^{K}$, we sample $K$ axial magnification vectors $\{\boldsymbol{\alpha}^{k}\}_{k=1}^{K}$ and a single degree of angle $\phi$. Then, we perform the same procedure as the next layers but with the axially magnified translation parameters $\{\boldsymbol{\alpha}^{k}(\operatorname{proj}_{\mathbf{\textbf{p}^{\phi}}}{\textbf{d}^{k}};\operatorname{proj}_{\mathbf{\textbf{p}^{\phi_{\perp}}}}{\textbf{d}^{k}})\}_{k=1}^{K}$. These previous, next, and axially magnified layer images and masks are then superimposed into a single image to yield $\textbf{I}_{1}$, $\textbf{I}_{2}$, and $\hat{\textbf{I}}^{\phi}$, respectively. Our dataset also includes the angle $\phi$ and the object-wise magnification map $\boldsymbol{\Lambda}$ which is generated by superimposing $\{\boldsymbol{\alpha}^{k}\}_{k=1}^{K}$ segmented with $\{\boldsymbol{\Omega}_{1}^{k}\}_{k=1}^{K}$. We observe that $\phi$ and $\boldsymbol{\Lambda}$ are useful to learn the representations robust to the noise, which will be discussed on Sec. 4.3. We provide more details in the supplementary materials.
Loss Function
Based on the implementation of DMM [23], we apply the reconstruction, texture, and shape losses. The reconstruction loss $L_{\text{recon}}$ is $l_{1}$ loss between the ground truth of the axially magnified image $\hat{\textbf{I}}^{\phi}$ and the estimated axially magnified image $\tilde{\textbf{I}}^{\phi}$. The texture loss $L_{\text{texture}}$ and the shape loss $L_{\text{shape}}$ are $l_{1}$ losses and designed so that the texture / shape representation represents intensity / motion information, respectively. We additionally apply perceptual loss $L_{\text{per}}$ [15] and Laplacian loss $L_{\text{lap}}$ between $\hat{\textbf{I}}^{\phi}$ and $\tilde{\textbf{I}}^{\phi}$. We observed an improvement in both quantitative and qualitative results when applying these two losses together. The total loss $L_{\text{total}}$ is as follows:
$L_{\text{total}}$ $=L_{\text{recon}}(\hat{\textbf{I}}^{\phi},\tilde{\textbf{I}}^{\phi})+\gamma L_{\text{lap}}(\hat{\textbf{I}}^{\phi},\tilde{\textbf{I}}^{\phi})+\beta\Bigl{(}L_{\text{texture}}(\textbf{T}_{1},\textbf{T}_{2})$
${\,+}L_{\text{shape}}(\textbf{S}^{x}_{2},\acute{\textbf{S}}^{x}_{2}){\,+}L_{\text{shape}}(\textbf{S}^{y}_{2},\acute{\textbf{S}}^{y}_{2}){\,+}L_{\text{per}}(\hat{\textbf{I}}^{\phi},\tilde{\textbf{I}}^{\phi})\Bigr{)},$ (7)
where $\acute{\textbf{S}}^{x}_{2}$ and $\acute{\textbf{S}}^{y}_{2}$ denote the shape representations of color perturbed next image $\acute{\textbf{I}}_{2}$ as suggested by Oh et al. [23]. We set $\beta{=}0.1$ and $\gamma{=}0.5$. Refer to the ablation study and the details of the loss functions in the supplementary material.
@@ 4 Experiments
We examine the performance of our method in comparison to other motion magnification methods, considering both cases of axial and generic motion magnification. We train our learning-based axial motion magnification network on the newly proposed dataset, which contains a total of 100k samples, for 50 epochs with a batch size of $8$ and a learning rate $2\times 10^{-4}$. We use both the dynamic and static modes in the experiments following DMM [23], without a temporal bandpass filter that separates the motion with the frequency of interest. Details and additional experiments can be found in the supplementary material, including the magnified results with the temporal bandpass filters.
@@ 4.1 Axial Motion Magnification
We evaluate our method compared to the phase-based method [37] in the axial motion magnification scenario, to demonstrate the effectiveness of the learning-based axial motion magnification. In our experiment, we modulate the phase-based to operate in axial scenario by employing a half-octave bandwidth pyramid and two orientations, with one of them having its phase representation manipulated along the axis of interest.
Figure 5: Quantitative results in axial motion magnification scenario. (a) In the subpixel test, our method demonstrates superior performance on SSIM over the phase-based method [37] across all input motion amount, ranging from $0.04$ to $1.0$. (b) In the noise tests when the input motion amount is 0.05 pixel, we observe a growing disparity in SSIM scores between our method and the phase-based approach, as the noise factor rises.
@@
We will demonstrate the advantage of our method that it can amplify only the motion along the axis of interest while disentangling the motions in uninterested directions that interfere with motion analysis. To illustrate this concept concretely, consider a scenario where a shaft is rotating in the radial direction. In such cases, magnifying and examining the motion along the axial direction, which is crucial to assess the condition of the rotating machinery [21], becomes challenging due to the dominance of rotational motion over the axial component. We conduct an experiment as depicted in Fig. 4 to illustrate this challenge, by attaching weights to a rotor to impose an imbalance, which results in axial vibrations. Then, we acquire a video of the imbalanced rotor, called rotor imbalance sequence, using a camera that captures $120$ Frames Per Second (FPS). We choose a horizontal-axis line in the original frame and plot $x$-t graphs for the magnified output frames from each method, respectively. Note that we also provide the result of Deep Motion Magnification (DMM) [23] as a reference to compare the results of axial motion magnification with generic motion magnification.
Figure 6: Motion legibility improvement. We visualize the 40$\times$ magnified frames of the structure, which are overlaid with the sampled trajectories from the KLT tracker. Ours shows simplified and legible motion trajectories when magnifying along the specific axis (i.e., y-axis in this case), while DMM [23] produces the trajectories that are more complex and challenging to interpret.
As shown in Fig. 4, our method produces the magnified output frames without artifacts and exhibits the $x$-t graph that clearly depicts axial vibrations. In contrast, the phase-based method [37] suffers from severe ringing artifacts, likely due to the overcompleteness of the complex steerable filter [28, 27], which cannot perfectly separate the phase representation into two orthogonal directions. DMM yields the magnified frames with artifacts and unclear axial vibrations in the $x$-t graph, since the representation of generic motion magnification method struggles to disentangle the dominant motion of the radial direction from the motion of interest, i.e., axial direction’s motion.
Figure 7: Qualitative results in generic motion magnification scenario. We amplify the motions in baby sequence with the magnification factor $\alpha=20$, using phase-based method [37], Singh et al. [29], STB-VMM [16], DMM [23], and our method. Both Ours and DMM favorably preserve the edges of the baby’s clothing and show no ringing artifacts in the magnified output frames and the $x$-t graphs. In contrast, the magnified output frames of the phase-based, STB-VMM, and Singh et al. show severe ringing artifacts or blurry results.
Figure 8: Quantitative results in generic motion magnification scenario. (a) In the subpixel test, which measures performance with varying input motion amount, Ours outperforms phase-based method [37] and Singh et al[29], and achieves favorable performance on SSIM compared to DMM [23] and STB-VMM [16]. (b) In the noise tests, compared to other methods, ours shows outstanding noise tolerance as the noise factor increases.
@@
To quantitatively evaluate our learning-based axial motion magnification method, we generate axial evaluation dataset based on the validation dataset of DMM [23]. The method of generating the dataset is almost the same as that of the training dataset. One difference is that we adjust the motion amplification factor to ensure that the amplified motion magnitude along a random axis is equal to 10. The motion amplification factor for the other axis is set to half the value. Note that we set $\phi$ to be 00 for this quantitative evaluation. We report the Structural Similarity Index (SSIM) [39] between the ground truth and output frames of the phase-based method and ours. As a reference, we provide the SSIM between ground truth and input frames. Figure 5 summarizes the results. We measure the SSIM by varying the levels of motion (Fig. 5-(a) Subpixel test) and additive noise (Fig. 5-(b) Noise test) in the input images. The number of evaluation data samples for each level of motion and noise is $1,000$. Regardless of the input motion magnitude and noise level, our method consistently outperforms the phase-based approach, which indicates that our proposed network architecture and dataset are effective for learning axis-wise disentangled representations.
@@
To demonstrate the efficacy of our method in improving the legibility of magnified motions, we use a structure that exhibits complex movements. We then visualize and compare the motion trajectories, tracked by the KLT tracker, of the 40×40$\times$40 × magnified video sequences of this structure using both the generic method (DMM) [23] and the axial method (Ours). As shown in Fig. 6, our method shows simplified and legible trajectories when magnifying along the specific axis (i.e., $y$-axis in this case), while DMM shows the trajectories difficult to decipher the motion of the structure.
@@ 4.2 Generic Motion Magnification
While our method focuses on the axial motion magnification task, it can be readily adapted for generic motion magnification scenario. This adaptability is achieved by simply multiplying the same magnification factors with the axis-wise shape representations. In the context of generic motion magnification, to demonstrate the effectiveness of our method, we compare it with the phase-based method [37] and the learning-based methods [23, 16, 29].
@@
We visualize the magnified output frames and plot the $x$-t graphs for the baby sequence, comparing ours with the several motion magnification methods in the generic motion magnification scenario (see Fig. LABEL:figure:generic_qualitative). Both our method and DMM [23] favorably preserve the edges of the baby’s clothing and show no ringing artifacts in the magnified results of breathing motion. In contrast, the phase-based method [37], STB-VMM [16], and Singh et al[29] show severe ringing artifacts or blurry results.
@@
To quantitatively verify the ability of our method in generic motion magnification, we synthesize the generic validation dataset. Unlike the axial motion magnification case, we set the motion magnification factor $\alpha$ to be identical along the $x$ and $y$ axes. We report SSIM [39] between ground truth frames and output frames from the phase-based method [37] and the learning-based methods [23, 29, 16]. As shown in Fig. 8, when the amount of input motion ranges from $0.04$ to $1.0$, ours outperforms the phase-based method and Singh et al[29], achieving favorable performance in SSIM compared to DMM [23] and STB-VMM [16]. Our method also shows notable noise tolerance as the noise factor rises in the noise test at $0.05$ pixel.
@@ 4.3 Ablation Study
In this section, we conduct ablation studies to evaluate the impact of the Motion Separation Module (MSM) and the components of the proposed synthetic training data. For the ablation studies, we carry out quantitative experiments on the evaluation dataset of both the generic case and the axial case that has random angles.
Figure 9: [Top] Architectural difference on the shape branch, [Bottom] Quantitative results of ablating Motion Separation Module (MSM). The modified DMM, desinged for ablation study, employs 2D convolutions and splits features along channel dimensions for axial motion magnification. Ours with Motion Separation Module (MSM) achieves higher SSIM in both the subpixel and noise tests on generic and axial evaluation datasets, respectively.
Motion Separation Module
To validate the effectiveness of our proposed Motion Separation Module (MSM), we design a competitor called modified DMM. While the design of the modified DMM closely resembles that of DMM [23], it differs in its approach to splitting the shape representation axis-wise for axial motion magnification. As shown in the top of Fig. 9, different from our method that uses 1D convolutions, the modified DMM employs 2D convolutions in the Shape branch and the Manipulator. The axis-wise shape representations of modified DMM are acquired by dividing the feature map along the channel dimension. We train the networks with the loss used in DMM [23] for an independent ablation on the network itself.
The bottom of Fig. 9 summarizes the quantitative results of our method and modified DMM. Ours with the MSM achieves higher SSIM on the subpixel and noise tests, compared to the modified DMM in both the axial and generic cases. These results show the effectiveness of the MSM in capturing small motions and handling noisy input images.
@@
To evaluate the impact of the angle $\phi$ and the object-wise motion magnification map $\boldsymbol{{\Lambda}}$, we generate the different types of training data varying the presence of these components. Our newly proposed dataset includes both $\phi$ and $\boldsymbol{{\Lambda}}$, whereas the dataset under the same configuration as DMM [23] lacks both. We further produce two additional datasets by individually adding each component to the dataset without $\phi$ and $\boldsymbol{{\Lambda}}$. Note that evaluating on the axial evaluation dataset is infeasible since it requires $\phi$ in the training time. As shown in Fig. 10, adding $\phi$ achieves no performance improvement in the subpixel test, making the network vulnerable to the noise of input images. In contrast, adding $\boldsymbol{{\Lambda}}$ leads to an increase on SSIM [39] in the subpixel test and huge robustness against noises in the noise test. The combined use of both $\phi$ and $\boldsymbol{{\Lambda}}$ yields the most significant performance improvement in the noise test, demonstrating that our proposed dataset is beneficial in the generic motion magnification task, as well.
Figure 10: Ablation study of the components in data generation pipeline in the generic evaluation dataset. We generate the different types of training data varying the presence of the angle $\phi$ and the object-wise motion magnification map $\boldsymbol{{\Lambda}}$, evaluating the networks trained on each dataset configuration using the generic evaluation dataset. Using both $\phi$ and $\boldsymbol{{\Lambda}}$ demonstrates superior noise tolerance compared to the other configurations.

### v3

@@ 3.3 Neural Networks and Training
Figure 3 shows the synthetic data generation pipeline. We sample one background from COCO [19] and $K{-}1$ number of foreground textures with segmentation masks from PASCAL VOC [10]. These elements are randomly located on image planes of resolution 384$\times$384 to produce $K$ previous layer images $\{{\mathbf{L}}^{k}_{1}\}_{k=1}^{K}$ and corresponding masks $\{\boldsymbol{\Omega}^{k}_{1}\}_{k=1}^{K}$. Following this, with randomly sampled $K$ translation parameters $\{\textbf{d}^{k}\}_{k=1}^{K}$, we generate the next layer images $\{{\mathbf{L}}^{k}_{2}\}_{k=1}^{K}$ and masks $\{\boldsymbol{\Omega}^{k}_{2}\}_{k=1}^{K}$ by translating the initial layers and masks according to $\{\textbf{d}^{k}\}_{k=1}^{K}$. For the axially magnified layer images $\{\hat{{\mathbf{L}}}^{\phi,k}\}_{k=1}^{K}$ and their masks $\{\hat{\boldsymbol{\Omega}}^{\phi,k}\}_{k=1}^{K}$, we sample $K$ axial magnification vectors $\{\boldsymbol{\alpha}^{k}\}_{k=1}^{K}$ and a single degree of angle $\phi$. Then, we perform the same procedure as the next layers but with the axially magnified translation parameters $\{\boldsymbol{\alpha}^{k}(\operatorname{proj}_{\mathbf{\textbf{p}^{\phi}}}{\textbf{d}^{k}};\operatorname{proj}_{\mathbf{\textbf{p}^{\phi_{\perp}}}}{\textbf{d}^{k}})\}_{k=1}^{K}$. These previous, next, and axially magnified layer images and masks are then superimposed into a single image to yield $\textbf{I}_{1}$, $\textbf{I}_{2}$, and $\hat{\textbf{I}}^{\phi}$, respectively. Our dataset also includes the angle $\phi$ and the object-wise magnification map $\boldsymbol{\Lambda}$ which is generated by superimposing $\{\boldsymbol{\alpha}^{k}\}_{k=1}^{K}$ segmented with $\{\boldsymbol{\Omega}_{1}^{k}\}_{k=1}^{K}$. We observe that utilizing both $\phi$ and $\boldsymbol{\Lambda}$ are useful for learning the representations distinguishing small motions from noises, which will be discussed on Sec. 4.3. Additionally, the adaptation of both $\phi$ and $\boldsymbol{\Lambda}$ enables pixel-wise axial motion magnification. We provide more details in the supplementary materials.
@@ 4 Experiments
Implementation Details
We train our learning-based axial motion magnification network on the newly proposed dataset, which contains a total of 100k samples, for 50 epochs with a batch size of $8$ and a learning rate $2\times 10^{-4}$.
Evaluation Setup
We examine the performance of our method in axial and generic motion magnification, respectively. In generic motion magnification, we compare our method to the phase-based method [39], Singh et al[31], STB-VMM [17], Pan et al[26], and DMM [24]. In axial motion magnification, there is no method of handling a user-specified angle and performing axial magnification due to our novel problem setup. Therefore, we propose a new axial baseline, called modified phase-based, by modifying Wadhwa et al. [39]. Specifically, we modulate the phase-based to operate in axial scenario by employing a half-octave bandwidth pyramid and two orientations, with one of them having its phase representation manipulated along the axis of interest. We use both the dynamic and static modes in the experiments following DMM [24]. Additional experiments of diverse scenarios and implementation details can be found in the supplementary material and video, including the magnified results with the temporal bandpass filters separating the motion with the frequency of interest.
@@ 4.1 Axial Motion Magnification
We evaluate our method compared to the modified phase-based method in the axial motion magnification scenario, to demonstrate the effectiveness of the learning-based axial motion magnification.
Figure 4: [Left] Imposing an imbalance on a rotor, [Right] Qualitative results in axial motion magnification scenario. We attach weights to a rotor to impose an imbalance and acquire rotor imbalance sequence, which has axial vibrations. Then, we amplify only the motion of rotor’s axial direction with the magnification factor $\alpha=40$, using ours and modified phase-based method. We also show the magnified result of DMM [24] as a reference result of generic motion magnification. Our method generates magnified frames without artifacts and exhibits the $x$-t slice showing clearly legible axial vibrations, while modified phase-based method and DMM both suffer from severe artifacts and have unclear axial vibrations in the $x$-t slice.
@@
We demonstrate the advantage of our method that it can amplify only the motion along the axis of interest while disentangling the motions in uninterested directions that interfere with motion analysis. To illustrate this concept concretely, consider a scenario where a shaft is rotating in the radial direction. In such cases, magnifying and examining the motion along the axial direction, which is crucial to assess the condition of the rotating machinery [22], becomes challenging due to the dominance of rotational motion over the axial component. We conduct an experiment shown in Fig. 4 by attaching weights to a rotor to impose an imbalance, which results in axial vibrations. Then, we acquire a video of the imbalanced rotor, called rotor imbalance sequence. We choose a horizontal-axis line in the original frame and visualize $x$-t slices for the magnified output frames from each method, respectively. Note that we also provide the result of DMM [24] as a reference to compare the results of axial motion magnification with generic motion magnification. As shown in Fig. 4, our method produces the magnified output frames without artifacts and exhibits the $x$-t slice that clearly depicts axial vibrations. In contrast, the modified phase-based method suffers from severe ringing artifacts, likely due to the overcompleteness of the complex steerable filter [30, 29], which cannot perfectly separate the phase representation into two orthogonal directions. DMM yields the magnified frames with artifacts and unclear axial vibrations in the $x$-t slice, since the representation of generic motion magnification method struggles to disentangle the dominant motion of the radial direction from the motion of interest, i.e., axial direction’s motion.
@@
Figure 5: Quantitative results in axial motion magnification scenario. (a) In the subpixel test, ours shows superior performance on SSIM over the modified phase-based method across all input motion amount, ranging from $0.04$ to $1.0$. (b) In the noise tests when the input motion amount is 0.05 pixel, we observe a growing disparity in SSIM scores between ours and the phase-based approach, as the noise factor rises.
To quantitatively evaluate our learning-based axial motion magnification method, we generate an axial evaluation dataset based on the validation dataset of DMM [24]. The method of generating the dataset is almost the same as that of the training dataset. One difference is that we adjust the motion amplification factor to ensure that the amplified motion magnitude along a random axis is equal to 10. The motion amplification factor for the other axis is set to half the value. Note that we set $\phi$ to be 00 for this quantitative evaluation. We report the Structural Similarity Index (SSIM) [41] between the ground truth and output frames of the modified phase-based method and ours. As a reference, we provide the SSIM between ground truth and input frames. Figure 5 summarizes the results. We measure the SSIM by varying the levels of motion (Fig. 5-(a) Subpixel test) and additive noise (Fig. 5-(b) Noise test) in the input images. The number of evaluation data samples for each level of motion and noise is $1,000$. Regardless of the input motion magnitude and noise level, our method consistently outperforms the modified phase-based approach, which indicates that our proposed network architecture and dataset are effective for learning axis-wise disentangled representations.
@@
Figure 6: Motion legibility improvement. We visualize the 40$\times$ magnified frames of the structure, which are overlaid with the sampled trajectories from the KLT tracker. Ours shows simplified and legible motion trajectories when magnifying along the specific axis (i.e., $x$-axis, $y$-axis or diagonal-axis in this case), while DMM [24] produces the trajectories that are more complex and hard to interpret.
To demonstrate the improved legibility of magnified motions by our method, we use a structure that exhibits complex movements. We then visualize and compare the motion trajectories, tracked by the KLT tracker, of the 40×40$\times$40 × magnified video sequences of this structure using both the generic method (DMM) [24] and the axial method (Ours). As shown in Fig. 6, our method shows legible trajectories when magnifying along the specific axis (i.e., $x$-axis, $y$-axis or diagonal-axis in this case), while DMM shows the entangled trajectories difficult to judge major motion characteristics.
@@ 4.2 Generic Motion Magnification
Our method can be readily adapted for generic motion magnification scenarios without further training. This adaptability is achieved by simply multiplying the same magnification factors with the axis-wise shape representations. In the context of generic motion magnification, we compare our method with the phase-based method [39] and the learning-based methods [24, 17, 31, 26].
Figure 7: Qualitative results in generic motion magnification scenario. We amplify the baby sequence with the magnification factor $\alpha{=}20$, using phase-based method [39], learning-based methods [31, 17, 24, 26], and Ours. Ours and DMM favorably preserve the edges of the clothes and show no ringing artifacts in the magnified frames and the $x$-t slices. In contrast, the magnified output frames of the phase-based, Singh et al., STB-VMM, and Pan et al. show ringing artifacts or blurry results.
Figure 8: Quantitative results in generic motion magnification scenario. (a) In the subpixel test, Ours outperforms phase-based method, Singh et al., and Pan et al. and achieves favorable performance on SSIM compared to DMM and STB-VMM. (b) In the noise test, Ours shows comparable noise tolerance compared to other methods and high noise tolerance as the noise factor increases.
@@
We visualize the magnified output frames and plot the $x$-t slices for the baby sequence, comparing ours with the several motion magnification methods in the generic scenarios (see Fig. 7). Both our method and DMM [24] favorably preserve the edges of the baby’s clothing and show no ringing artifacts in the magnified results of breathing motion. In contrast, the phase-based method [39], Singh et al[31], STB-VMM [17], Pan et al[26] and show severe ringing artifacts or blurry results [^1] We reproduced all the results using the codes publicly accessible..
@@
To quantitatively verify the ability of our method in generic motion magnification, we synthesize a generic validation dataset. Unlike the axial case, we set the magnification factor $\alpha$ to be identical along the $x$ and $y$ axes. As shown in Fig. 8, we report SSIM [41] between ground truth and output frames from the phase-based method [39] and the learning-based methods [24, 31, 17, 26]. For input motion ranges from $0.04$ to $1.0$, ours outperforms the phase-based method, Singh et al[31], Pan et al[26]. Compared to DMM [24] and STB-VMM [17], ours demonstrates favorable performance, which exceeds the threshold for visually acceptable SSIM scores [15]. Ours demonstrates comparable noise tolerance to other methods and exhibits high noise tolerance as noise factor increases.
@@ 4.3 Ablation Study
In this section, we conduct ablation studies to evaluate the impact of the Motion Separation Module (MSM) and the components of the proposed synthetic training data. We carry out quantitative experiments on the evaluation dataset of both the generic case and the axial case that has random angles.
Motion Separation Module (MSM)
Figure 9: [Top] Architectural difference on the shape branch, [Bottom] Quantitative results of ablating Motion Separation Module (MSM). (a) The modified DMM, designed for ablation study, employs 2D convolutions and splits features along channel dimensions for axial motion magnification. (c) Ours with MSM generally achieves higher SSIM in the subpixel test on generic and axial evaluation datasets. (d) In the noise test, Ours shows comparable performance to the modified DMM.
To validate the effectiveness of MSM, we design a competitor called modified DMM, which closely resembles that of DMM [24]. As shown in the top of Fig. 9, different from our method that uses 1D convolutions, the modified DMM employs 2D convolutions in the Shape branch and the Manipulator. The axial shape representations of the modified DMM are acquired by dividing the feature map along the channel dimension. We train the networks with the same loss function and training details as Ours. The bottom of Fig. 9 shows that Ours with MSM generally achieves higher SSIM in the subpixel test on the generic and axial evaluation datasets. These results show the effectiveness of the MSM in capturing small motions. In the noise test, Ours shows comparable performance to the modified DMM.
Figure 10: Ablation study of the components in data generation in the generic evaluation dataset. We generate the different training data varying the presence of the angle $\phi$ and the object-wise motion magnification map $\boldsymbol{{\Lambda}}$, and evaluate the networks trained on each dataset configuration using the generic evaluation dataset. (a) Using both $\phi$ and $\boldsymbol{{\Lambda}}$ demonstrates best performance in the subpixel test. (b) In the noise test, we observe that utilizing $\boldsymbol{{\Lambda}}$ notably enhances noise tolerance.
@@
To evaluate the impact of the angle $\phi$ and the object-wise motion magnification map $\boldsymbol{{\Lambda}}$, we generate the different types of training data varying the presence of these components. Our newly designed dataset incorporates both $\phi$ and $\boldsymbol{\Lambda}$, contrasting with the dataset that follows the same setup as DMM [24], which does not contain either element. In addition, we generate two more datasets that each add one of these components (i.e., either $\phi$ or $\boldsymbol{\Lambda}$) to the base dataset that initially does not include them. Note that evaluating the networks trained on these datasets on the axial evaluation dataset is infeasible since the networks trained without $\phi$ cannot perform axial motion magnification. Thus, we use the generic evaluation dataset for this ablation study. Fig. 10 shows that the addition of either $\phi$ or $\boldsymbol{{\Lambda}}$ achieves no improvement in the subpixel test. The combined use of both $\phi$ and $\boldsymbol{{\Lambda}}$ yields the most significant performance improvement in the subpixel test, demonstrating that our proposed data set is beneficial in the generic motion magnification task as well. In the noise test, utilizing $\boldsymbol{{\Lambda}}$ notably enhances noise tolerance, while the addition of $\phi$ has no effect on noise tolerance.